I am stuck, yet again :)
Express the quadratic function p(x)=a(x-h)^2+k in standard form, and identify a, b and c.
Thanks,in advance, for your help

Answer by stanbon(75887)

Express the quadratic function p(x)=a(x-h)^2+k
in standard form, and identify a, b and c.
-------
p(x) = a[x^2 - 2hx + h^2] + k
-----
p(x) = ax^2 - 2ahx + ah^2 + k
-----
a = a
b = -2ab
c = ah^2+k
------------------
Cheers,
Stan H.
--------------